# Rankers, Judges, and Assistants: Towards Understanding the Interplay of LLMs in Information Retrieval Evaluation

### v1

@@ 5.2 Open Questions and Future Directions
  • Assessing and Improving LLM Judge Quality. Our findings highlight the critical importance of LLM judge quality, revealing shortcomings in their discriminative ability and biases toward LLM-powered rankers. Developing robust methods for assessing and improving the quality of LLM judges is a crucial research direction for the IR community, potentially drawing motivations from horizontal autorater efforts [57].
  • Robustness Against Adversarial Attacks. LLM judges, similar to LLM-based rankers, are susceptible to adversarial attacks, including keyword stuffing and content injection [2, 52]. Understanding these vulnerabilities and developing effective mechanism to enhance the robustness of LLM judges against them are critical areas for ensuring the practical applicability of LLM judges in real-world scenarios.
  • Human-in-the-Loop LLM Judges. The potential for using LLMs to augment human assessors, e.g., for quality-control, has been suggested [13, 48]. While Upadhyay et al. [55] found that human-in-the-loop processes did not bring obvious tangible benefits, their study represents a preliminary investigation. Further research is needed to explore the potential of this approach more comprehensively.
  • From Passages to Longer Documents. Most of the existing work focuses on paragraphs as the unit of retrieval, using either the TREC DL [30, 13, 2, 5] or RAG [55] benchmarks. There is much less work on ad hoc retrieval, with exceptions including TREC-8 [13] and Robust [54]. It is known that LLMs handle long context differently [23] and its implication in judging long documents need further investigation.
@@
  • Domain-specific Solutions. While most existing research focuses on general-purpose search, specific domains might require purpose-built solutions. Recent work, for example, has explored the use of LLM judges for e-commerce search [32, 46]. Extending this concept of domain specialization, the applicability and value of LLM judges in specialized domains requiring expert knowledge (e.g., medical or legal search) remain less clear. In such domains, LLM judges might offer potential cost savings and more in-depth domain-specific knowledge compared to non-expert human assessors, but they also introduce new challenges, including the need for high accuracy, the potential for serious consequences from errors, and the complexities of expert judgment. Future research should explore the use of LLM judges in these contexts, carefully considering the trade-offs between cost, efficiency, and the risks of inaccurate assessments.
@@
  • Training Models on LLM-generated Labels. Training retrieval models on data labeled by LLMs introduces a significant risk of circularity and bias amplification. If done recursively, this might lead to model collapse [47]. Thomas et al. [54] acknowledge that parts of the Bing search engine are retrained using LLM-generated labels. Understanding the long-term effects of such training is an important research direction.
6 Conclusion

### v2

@@ 5.2. Open Questions and Future Directions
  • Assessing and Improving LLM Judge Quality. Our findings highlight the critical importance of LLM judge quality, revealing shortcomings in their discriminative ability and biases toward LLM-powered rankers. Developing robust methods for assessing and improving the quality of LLM judges is a crucial research direction for the IR community, potentially drawing motivations from horizontal autorater efforts (Vu et al., 2024).
  • Robustness Against Adversarial Attacks. LLM judges, similar to LLM-based rankers, are susceptible to adversarial attacks, including keyword stuffing and content injection (Alaofi et al., 2024; Tamber and Lin, 2025). Understanding these vulnerabilities and developing effective mechanism to enhance the robustness of LLM judges against them are critical areas for ensuring the practical applicability of LLM judges in real-world scenarios.
  • Human-in-the-Loop LLM Judges. The potential for using LLMs to augment human assessors, e.g., for quality-control, has been suggested (Faggioli et al., 2023; Soboroff, 2025). While Upadhyay et al. (2024a) found that human-in-the-loop processes did not bring obvious tangible benefits, their study represents a preliminary investigation. Further research is needed to explore the potential of this approach more comprehensively.
  • From Passages to Longer Documents. Most of the existing work focuses on paragraphs as the unit of retrieval, using either the TREC DL (MacAvaney and Soldaini, 2023; Faggioli et al., 2023; Alaofi et al., 2024; Chen et al., 2024a) or RAG (Upadhyay et al., 2024a) benchmarks. There is much less work on ad hoc retrieval, with exceptions including TREC-8 (Faggioli et al., 2023) and Robust (Thomas et al., 2024). It is known that LLMs handle long context differently (Liu et al., 2024a) and its implication in judging long documents need further investigation.
@@
  • Domain-specific Solutions. While most existing research focuses on general-purpose search, specific domains might require purpose-built solutions. Recent work, for example, has explored the use of LLM judges for e-commerce search (Mehrdad et al., 2024; Sachdev et al., 2025). Extending this concept of domain specialization, the applicability and value of LLM judges in specialized domains requiring expert knowledge (e.g., medical or legal search) remain less clear. In such domains, LLM judges might offer potential cost savings and more in-depth domain-specific knowledge compared to non-expert human assessors, but they also introduce new challenges, including the need for high accuracy, the potential for serious consequences from errors, and the complexities of expert judgment. Future research should explore the use of LLM judges in these contexts, carefully considering the trade-offs between cost, efficiency, and the risks of inaccurate assessments.
@@
  • Training Models on LLM-generated Labels. Training retrieval models on data labeled by LLMs introduces a significant risk of circularity and bias amplification. If done recursively, this might lead to model collapse (Shumailov et al., 2024). Thomas et al. (2024) acknowledge that parts of the Bing search engine are retrained using LLM-generated labels. Understanding the long-term effects of such training is an important research direction.
6. Conclusion